kcn has a density of 1.52g/cm3 and crystallines in nacl type structure calculate the distance between k+ and cn- ions

Grade:12

1 Answers

Arun
6 years ago
Density D = z* M/ a3 * Na
z = number of atoms per unit cell
KCN is NaCl type structure hence z =4
M = molar mass of KCN = 65.12 gm/mol
a = edge length
Na = 6.023 * 1023
putting all the value in formula
 
a3 = 4 *65.12/ 1.52 * 6.023 * 1023
a3 = 330.75 * 10 –24
a = 6.915 * 10 –8 cm
Edge length in rock salt (a) = 2(rc + ra)
(rc + ra) = a/ 2 = 3.4575 * 10 –8
so distance will be 3.4575 * 10 –8 cm or 345.75 pico metre